40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3165     Entry time: Wed Jul 7 11:23:08 2010
Author: Sharmila, Katharine 
Type: Update 
Category: WIKI-40M Update 
Subject:  

Weekly Update:

Last Weds-Thurs, we wrote and edited our progress reports.

Tuesday (and Weds morning):  Continued circuit analysis of Haixing's circuit and plotting transfer functions (almost have one for entire circuit).  Hooked up OSEM and circuit to power supply, but the LED didn't appear to light up in IR.  Now we are going to hook the OSEM directly to the power supply, sans circuit, to see if the problem is with the circuit or OSEM.

ELOG V3.1.3-